An electric bell takes a current of 0.2 ampere from a battery of 2 dry cells connected in a series each cell has electric motive force of 1.5 volt and an internal resistance of 1.0 ohms . calculate the effective resistive of the cell
1
Expert's answer
2020-09-30T12:25:54-0400

solution

current in bell (I)=0.2 A

EMF(E)=1.5 volt

internal resistance(r)=1.0 ohm

total resistance

(Rt)=Rbell + 2r


(Rt)=Rbell + 2"\cdot 1"


Rbell = Rt - 2


total resistance can be calculated as

"R_t=\frac{2E}{I}"


"R_t=\frac{2\times1.5}{0.2}\\=15\Omega"


then

"R_{bell}=15-2\\=13\Omega"


so resistance of bell is 13 ohm .